OpenStudy (anonymous):

here's one that seems a little bit tough: Test it's convergence (absolutely, conditionally, or divergent) \[\sum_{n=1}^{\infty}{(-1)^n\frac{2^n*n!}{5*8*11*...*(3n+2)}}\]

OpenStudy (anonymous):

its*

OpenStudy (zarkon):

use ratio test

OpenStudy (anonymous):

yeah that's what I thought.... but I hate the ratio test :-(

OpenStudy (anonymous):

integral test and root test are out of the question. Haven't tried the simple limit test; don't know if it applies here.

OpenStudy (zarkon):

use the ratio test...it really isn't that bad

OpenStudy (anonymous):

simple limit test gives me a negative infinity answer... so it is divergent :-D

OpenStudy (anonymous):

or am I mistaken? what does the ratio test tell me?

OpenStudy (zarkon):

looks like it converges to me

OpenStudy (anonymous):

the magic of computing!

OpenStudy (zarkon):

I don't think that is correct

OpenStudy (zarkon):

Mathematica version 8 says it converges

OpenStudy (anonymous):

I am willing to believe Mathematica 8, but does it have like a working out or a proof?

OpenStudy (zarkon):

formatting for the copy is a little messed up ... but In[2]:= SumConvergence[((-1)^n*2^n*Factorial[n])/(\!\( \*UnderoverscriptBox[\(\[Product]\), \(k = 1\), \(n\)]\(3 k\)\)+2),n] Out[2]= True

OpenStudy (zarkon):

the ratio test works...gives me 2/3 which is less than 1

OpenStudy (anonymous):

I better try it.... I'm just so scared of the 2^n and the 6*8*11*.....

OpenStudy (zarkon):

\[\left|\frac{a_{n+1}}{a_n}\right|=\left|\frac{(-1)^{n+1}\frac{2^{n+1}*(n+1)!}{5*8*11*...*(3n+2)(3n+5)}}{(-1)^n\frac{2^n*n!}{5*8*11*...*(3n+2)}}\right|\] \[=\frac{2^{n+1}*(n+1)!}{5*8*11*...*(3n+2)(3n+5)}{\frac{5*8*11*...*(3n+2)}{2^n*n!}}\] \[=\frac{2(n+1)}{3n+5}\to\frac{2}{3}<1\]

OpenStudy (anonymous):

doesn't look so bad... but why would wolfram be wrong :-(... I trusted him so much.

OpenStudy (zarkon):

not sure..since it is the same company that makes mathematica

OpenStudy (zarkon):

it's not the first time I've seen wolfram wrong

OpenStudy (anonymous):

what else does it do wrong?

OpenStudy (zarkon):

I don't remember the specific problem